# Crosstalk-Robust Quantum Control in Multimode Bosonic Systems

### v1

@@ III.1.2 Optimization procedure
Figure 4: Optimization results for the $\hat{X}_{\pi}$ gate. (a) Average infidelities in Eq. (19) versus pulse durations for both closed (dashed line) and open (solid line) systems. The optimal duration selected from the range considered for this optimization is 20 ns (indicated by a star). (b) $I$ and $Q$ quadratures of the pulse envelopes for the selected 20-ns pulse from (a). (c) The frequency components of the pulse envelopes, corresponding to the 20-ns pulse. (d) Comparative analysis of the 20-ns pulses, from both QOC (blue) and DRAG (red) pulses, in terms of their infidelities as a function of detuning. The average and standard deviation of the infidelities are tabulated correspondingly by considering a uniform distribution of $\rho(\delta)$ for simplicity. The result shows that the optimized pulse is more robust than the commonly used DRAG pulse. [Parameters: $\delta_{i}/2\pi\in\{\pm 1,\pm 3,\pm 4,\pm 7,\pm 10\}$ MHz, $K_{\text{q}}/2\pi=-200$ MHz, $T_{1}=50$ $\mu$s, $T_{\phi}=50$ $\mu$s]
Quantum optimal control has been used to explore robust control schemes by minimizing the cost function that encodes the closed- or open-system average infidelity [70, 71, 72, 73, 74]
@@ III.1.3 Benchmarking for an isolated transmon ancilla
We benchmark our optimization procedure for a transmon with typical parameters: $K_{\text{q}}/2\pi=-200$ MHz, $T_{1}=50$ $\mu$s, and $T_{\phi}=50$ $\mu$s. The optimization results for the $\hat{X}_{\pi}$ gate are shown in Fig. 4. Further details for the $\hat{X}_{\pi/2}$ gate are provided in Appendix C. Figure 4(a) shows the average infidelities obtained for closed- and open-system simulations with varying pulse durations. Without decoherence, extending the pulse duration generally leads to a reduction in the cost value. In contrast, in the presence of decoherence, utilizing shorter pulses is more effective in reducing the infidelity. Given the particular transmon parameters under consideration, a pulse duration of 20 ns achieves the lowest cost. The temporal profiles and frequency components of the optimized 20-ns pulse are shown in Fig. 4(b) and (c), respectively. In Fig. 4(d), we study the infidelity as a function of detuning, comparing the performance between the optimized and the DRAG pulses. In both closed- and open-system simulations, the optimized pulse consistently demonstrates substantially reduced infidelities across almost all detunings. Consequently, the optimized pulse exhibits enhanced robustness against frequency detuning in the transmon ancilla. The open-system infidelities obtained from these two pulses are limited by distinct factors: the infidelity derived from the QOC pulse is dominated by decoherence errors, while the one generated from the DRAG pulse is limited by coherent crosstalk. Moreover, we demonstrate that the optimized pulse we propose exhibits enhanced robustness compared to the one featured in Ref. 68, as detailed in Appendix C.
@@ III.2 Robust control of a composite cavity-ancilla system
Figure 5: Infidelities of $\hat{X}_{\pi}$ gate pulses acting on an ancilla dispersively coupled to a cavity mode. The horizontal axis denotes the dispersive shift $\chi$. Solid and dashed lines show the infidelities achieved with and without the cavity phase corrections in Eq. (29), respectively. For reference, infidelities obtained without relative cavity phase are displayed using discrete markers. [Parameters: $K_{\text{q}}/2\pi=-200\,\text{MHz}$, $d_{\text{c}}=10$]
@@
where $\hat{P}^{\prime}_{0,1}=\hat{P}_{0,1}\otimes\hat{\mathbb{I}}$ represents the projector onto the ancilla computational subspace, $\hat{U}_{\text{target}}^{\prime}=\hat{U}_{\text{target}}\otimes\hat{\mathbb{I}}$ describes the target operation of the full system, and $d_{\text{c}}$ denotes the cavity cutoff dimension. In Fig. 5, we present the ideal scenario where there is no relative phase associated with the cavity mode, i.e., $\theta(\delta)$ remains constant. This case is used as a reference and is depicted by circular dots. In practice, the infidelity, displayed with the dashed curve, shows a minor deviation from the ideal case. Remarkably, this deviation can be almost entirely eliminated using the correction method outlined in Eq. (29), with the outcomes displayed as the solid line. Hence, the numerical simulation corroborates the preceding analytical proof, affirming the applicability of a robust pulse optimized for a single ancilla to coupled systems, with minor modifications.
@@ IV.1 Quantum state transfer with an idling mode
Figure 6: Benchmarking of QOC and DRAG pulses in Fock state generation, $|0\rangle|\alpha\rangle|g\rangle\to|n\rangle|\alpha\rangle|g\rangle$. (a) Infidelity as a function of mean photon number $|\alpha|^{2}$ in the idling mode, for the preparation of target Fock states with $n=\{1,3,5\}$. DRAG pulse infidelity increases with $|\alpha|^{2}$, while QOC pulse infidelity remains nearly independent of $|\alpha|^{2}$. (b) Wigner tomography $\mathcal{W}(\eta)$ visualizing the state of the target mode subsystem for $|\alpha|^{2}=16$. The fidelities of states obtained using QOC and DRAG are 99.1% and 73.3%, respectively. [Parameters: $K_{\text{q}}/2\pi=-200$ MHz, $\chi_{1}/2\pi=\chi_{2}/2\pi=-300$ kHz]
@@
We compute the state-transfer infidelity of the composite system as $1-|\langle\psi_{\text{f}}|\psi_{\text{target}}|^{2}$, for the final state $|\psi_{\text{f}}\rangle$ and the target state $|\psi_{\text{target}}\rangle=|n\rangle|\alpha\rangle|g\rangle$, and plot its dependence on $|\alpha|^{2}$ in Fig. 6(a). When the idling mode is in the vacuum state (i.e., $\alpha=0$), the corresponding infidelities act as reference values, excluding crosstalk errors from the idling mode. As the average photon number in the idling mode increases, the infidelities associated with the DRAG scheme exhibit a marked escalation. Conversely, in the case of QOC, the infidelities barely deviate from the reference one, despite large variations in the average photon number of the idling mode. For instance, we consider the case of creating the Fock state $n=5$, when the idling mode has a mean photon number of $|\alpha|^{2}=16$. The infidelity of the final state generated by DRAG pulses surpasses $0.25$, whereas it remains below $0.01$ for QOC pulses. This marked difference illustrates the limited performance of conventional DRAG pulses in the presence of crosstalk, and underscores the superior robustness of the developed QOC pulses.
@@ IV.2 Two-mode entanglement operations
Figure 7: Benchmarking the robustness of QOC and DRAG pulses in the generation of entangled Bell-cat states, $|0\rangle|0\rangle|g\rangle\to\mathcal{N}(|\alpha\rangle|\alpha\rangle+|{-}\alpha\rangle|{-}\alpha\rangle)|g\rangle$, with $\mathcal{N}\approx 1/\sqrt{2}$. (a) Infidelity as a function of the mean photon number $|\alpha|^{2}$, with fixed dispersive shifts $|\chi_{1}|/2\pi=|\chi_{2}|/2\pi=200$ kHz. Both pulses have a duration of 20 ns. (b) and (c) Infidelity versus the dispersive shifts, with a constant $|\alpha|^{2}=16$, for ancilla pulse durations of 20 and 50 ns, respectively. Solid and dashed lines are obtained from simulations with and without decoherence errors. In all scenarios considered here, the infidelities derived from QOC pulses are significantly lower compared with the ones obtained from DRAG pulses. (d) and (f) Joint characteristic function of the Bell-cat state, evaluated at real $(\eta_{1},\eta_{2})$ and imaginary arguments $(i\eta_{1},i\eta_{2})$, respectively. The simulation is performed for an open system with 20-ns QOC pulses. (e) and (g) 1D line cuts obtained from (d) and (f). While the results derived from QOC pulses (blue) closely align with the ideal case (gray), those obtained from DRAG pulses (red) exhibit clear deviations. [Parameters: $K_{\text{q}}/2\pi=-200$ MHz, $T_{1}=T_{\phi}=50$$\mu$s, $T_{\phi 1,2}=150$ ms]
@@
V Conclusions
@@ Appendix A Multimode universal control with ECD protocol
Appendix B Ancilla gate decomposition for efficient experimental calibration

### v2

@@ III.1.2 Optimization procedure
Figure 4: Optimization results for the $\hat{X}_{\pi}$ gate. (a) Average infidelities in Eq. (12) versus pulse durations for both closed (dashed line) and open (solid line) systems. The optimal duration selected from the range considered for this optimization is 20 ns (indicated by a star). (b) $I$ and $Q$ quadratures of the pulse envelopes for the selected 20-ns pulse from (a). (c) The frequency components of the pulse envelopes, corresponding to the 20-ns pulse. (d) Comparative analysis of the 20-ns pulses, from both QOC (blue) and DRAG (red) pulses, in terms of their infidelities as a function of detuning. The average and standard deviation of the infidelities are tabulated correspondingly by considering a uniform distribution of $\rho(\delta)$ for simplicity. The result shows that the optimized pulse is more robust than the commonly used DRAG pulse. [Parameters: $\delta_{i}/2\pi\in\{\pm 1,\pm 3,\pm 4,\pm 7,\pm 10\}$ MHz, $K_{\text{q}}/2\pi=-200$ MHz, $T_{1}=50$ $\mu$s, $T_{\phi}=50$ $\mu$s.]
Quantum optimal control has been extensively used to develop robust control schemes by minimizing cost functions that encode robustness metrics. One approach is to penalize the derivative of the closed- or open-system fidelity with respect to detuning $\delta$ [70, 67, 68]. However, a more intuitive and pragmatic method is to penalize the average infidelity [71, 72, 73, 74, 75]
@@ III.1.3 Benchmarking for an isolated transmon ancilla
We benchmark our optimization procedure for a transmon with typical parameters: $K_{\text{q}}/2\pi=-200$ MHz, $T_{1}=50$ $\mu$s, and $T_{\phi}=50$ $\mu$s. The optimization results for the $\hat{X}_{\pi}$ gate are shown in Fig. 4. Further details for the $\hat{X}_{\pi/2}$ gate are provided in Appendix E. Figure 4(a) shows the average infidelities obtained for closed- and open-system simulations with varying pulse durations. Without decoherence, extending the pulse duration generally leads to a reduction in the cost value. In contrast, in the presence of decoherence, utilizing shorter pulses is more effective in reducing the infidelity. Given the particular transmon parameters under consideration, a pulse duration of 20 ns achieves the lowest cost. The temporal profiles and frequency components of the optimized 20-ns pulse are shown in Fig. 4(b) and (c), respectively. In Fig. 4(d), we study the infidelity as a function of detuning, comparing the performance between the optimized and the DRAG pulses. In both closed- and open-system simulations, the optimized pulse consistently demonstrates substantially reduced infidelities across almost all detunings. Consequently, the optimized pulse exhibits enhanced robustness against frequency detuning in the transmon ancilla. The open-system infidelities obtained from these two pulses are limited by distinct factors: the infidelity derived from the QOC pulse is dominated by decoherence errors, while the one generated from the DRAG pulse is limited by coherent crosstalk. Moreover, we demonstrate that the optimized pulse we propose exhibits enhanced robustness compared to the one featured in Ref. 67, as detailed in Appendix E.
@@ III.2 Robust control of a composite cavity-ancilla system
Figure 5: Infidelities of $\hat{X}_{\pi}$ gate pulses acting on an ancilla dispersively coupled to a cavity mode. The horizontal axis denotes the dispersive shift $\chi$. Solid and dashed lines show the infidelities achieved with and without the cavity phase corrections in Eq. (22), respectively. For reference, infidelities obtained without relative cavity phase are displayed using discrete markers. [Parameters: $K_{\text{q}}/2\pi=-200\,\text{MHz}$, $d_{\text{c}}=10$, with other nonlinearities derived from Eq. (26).]
@@
where $\hat{P}^{\prime}_{0,1}=\hat{P}_{0,1}\otimes\hat{\mathbb{I}}$ represents the projector onto the ancilla computational subspace, $\hat{U}_{\text{target}}^{\prime}=\hat{U}_{\text{target}}\otimes\hat{\mathbb{I}}$ describes the target operation of the full system, and $d_{\text{c}}$ denotes the cavity cutoff dimension. In Fig. 5, we present the ideal scenario where there is no relative phase associated with the cavity mode, i.e., $\theta(\delta)$ remains constant. This case is used as a reference and is depicted by circular dots. In practice, the infidelity, displayed with the dashed curve, shows a minor deviation from the ideal case. Remarkably, this deviation can be almost entirely eliminated using the correction method outlined in Eq. (22), with the outcomes displayed as the solid line. Hence, the numerical simulation corroborates the preceding analytical proof, affirming the applicability of a robust pulse optimized for a single ancilla to coupled systems, with minor modifications. It is crucial to emphasize that the virtual phase correction is only effective when the detuning remains within the range of robustness, making it primarily compatible with pulses obtained through optimal control techniques. In Appendix F, we elaborate on the limitations of this method, supported by examples involving Gaussian and DRAG pulses.
@@ IV.1 Quantum state transfer with an idling mode
Figure 6: Benchmarking of QOC and DRAG pulses in Fock state generation, $|0\rangle|\alpha\rangle|g\rangle\to|n\rangle|\alpha\rangle|g\rangle$. (a) Infidelity as a function of mean photon number $|\alpha|^{2}$ in the idling mode, for the preparation of target Fock states with $n=\{1,3,5\}$. DRAG pulse infidelity increases with $|\alpha|^{2}$, while QOC pulse infidelity remains nearly independent of $|\alpha|^{2}$. (b) Wigner tomography $\mathcal{W}(\eta)$ visualizing the state of the target mode subsystem for $|\alpha|^{2}=16$. The fidelities of states obtained using QOC and DRAG are 99.1% and 73.3%, respectively. [Parameters: $K_{\text{q}}/2\pi=-200$ MHz, $\chi_{1}/2\pi=\chi_{2}/2\pi=-300$ kHz, $\Delta_{1}/2\pi=\Delta_{2}/2\pi=2\,\text{GHz}$, with other nonlinearities derived from Eq. (26).]
@@
We compute the state-transfer infidelity of the composite system as $1-|\langle\psi_{\text{f}}|\psi_{\text{target}}\rangle|^{2}$, for the final state $|\psi_{\text{f}}\rangle$ and the target state $|\psi_{\text{target}}\rangle=|n\rangle|\alpha\rangle|g\rangle$, and plot its dependence on $|\alpha|^{2}$ in Fig. 6(a). When the idling mode is in the vacuum state (i.e., $\alpha=0$), the corresponding infidelities act as reference values, excluding crosstalk errors from the idling mode. As the average photon number in the idling mode increases, the infidelities associated with the DRAG scheme exhibit a marked escalation. Conversely, in the case of QOC, the infidelities barely deviate from the reference one, despite large variations in the average photon number of the idling mode. For instance, we consider the case of creating the Fock state $n=5$, when the idling mode has a mean photon number of $|\alpha|^{2}=16$. The infidelity of the final state generated by DRAG pulses surpasses $0.25$, whereas it remains below $0.01$ for QOC pulses. This marked difference illustrates the limited performance of conventional DRAG pulses in the presence of crosstalk, and underscores the superior robustness of the developed QOC pulses.
@@ IV.2 Two-mode entanglement operations
Figure 7: Benchmarking the robustness of QOC and DRAG pulses in the generation of entangled Bell-cat states, $|0\rangle|0\rangle|g\rangle\to\mathcal{N}(|\alpha\rangle|\alpha\rangle+|{-}\alpha\rangle|{-}\alpha\rangle)|g\rangle$, with $\mathcal{N}\approx 1/\sqrt{2}$. (a) Infidelity as a function of the mean photon number $|\alpha|^{2}$, with fixed dispersive shifts $|\chi_{1}|/2\pi=|\chi_{2}|/2\pi=200$ kHz. Both pulses have a duration of 20 ns. (b) and (c) Infidelity versus the dispersive shifts, with a constant $|\alpha|^{2}=16$, for ancilla pulse durations of 20 and 50 ns, respectively. Solid and dashed lines are obtained from simulations with and without decoherence errors. In all scenarios considered here, the infidelities derived from QOC pulses are significantly lower compared with the ones obtained from DRAG pulses. (d) and (f) Joint characteristic function of the Bell-cat state, evaluated at real $(\eta_{1},\eta_{2})$ and imaginary arguments $(i\eta_{1},i\eta_{2})$, respectively. The simulation is performed for an open system with 20-ns QOC pulses. (e) and (g) 1D line cuts obtained from (d) and (f). While the results derived from QOC pulses (blue) closely align with the ideal case (gray), those obtained from DRAG pulses (red) exhibit clear deviations. [Parameters: $K_{\text{q}}/2\pi=-200$ MHz, $\Delta_{1}/2\pi=\Delta_{2}/2\pi=2\,\text{GHz}$, $T_{1}=T_{\phi}=50$$\mu$s, $T_{\phi 1,2}=150$ ms, with other nonlinearities and decoherence parameters derived from Eq. (26) and Eq. (32), respectively.]
@@
V Discussion and outlook
The robust control protocol for two-mode ECD gates significantly suppresses crosstalk errors in “all-to-one” multimode cavity-based quantum processor architectures. However, neither the “all-to-one” architecture nor the ECD gate is the only viable option for controlling multimode oscillators. In this section, we first examine the advantages and challenges of the chosen architecture and compare it with other existing architectures in the literature. Subsequently, we discuss how our robust control technique can be applied to other multimode gate protocols, extending its benefits beyond the specific implementations discussed in this paper.
V.1 Comparison with other multimode architectures
In the “all-to-one” multimode architecture discussed in this work, a single ancilla is coupled to every cavity mode to execute single-mode gate operations. For multimode operations, the same ancilla entangles with multiple cavity modes to mediate the entanglement between modes that are otherwise not directly coupled. In addition to the two-mode ECD protocol presented here, this architecture supports various other multimode protocols that create multimode entanglement deterministically [86, 27, 37]. In general, this architecture features several advantages: First, the usage of a single transmon ancilla greatly reduces the effort in device fabrication and operation. Additionally, the architecture minimizes the number of lossy ancillas, which typically Purcell-limits the cavity coherence. Moreover, weakly coupled ancilla can be used to further suppress Purcell loss while still performing fast gates through various acceleration strategies [36, 37]. However, this architecture also presents challenges. The ancilla becomes entangled with the cavity modes during multimode operations, making ancilla decoherence the dominant source of error.
In an alternative architecture, each cavity mode is controlled by its own transmon ancilla, with a coupler facilitating the parametric beamsplitter interaction between cavity modes. For example, this can be realized with three-wave or four-wave mixing process in a Superconducting Nonlinear Asymmetric Inductive eLement (SNAIL) [87, 88], a DC Superconducting QUantum Interference Device (SQUID) [89], and a transmon coupler [90, 91]. When the coupler is undriven, different cavity modes are decoupled, eliminating crosstalk errors. This then allows the usage of selective single-mode control such as the SNAP gate. Additionally, the coupler is virtually populated during the beamsplitter interaction [90, 92], so the fidelity of multimode operations is less affected by coupler decoherence. However, this architecture demands a more complex experimental setup, requiring a transmon ancilla for each cavity mode in addition to the coupler. All the individual ancillas and the coupler introduce additional Purcell loss to the cavity modes. Moreover, strong coupling between the cavity modes and the coupler is typically required to enable fast inter-cavity operations, which further limits the cavity coherence.
The optimal architecture for multimode control remains an open question. A potential hybrid solution is the “manipulator-storage” architecture [53, 93, 94], which combines features of both strategies. It consists of a manipulator unit with a few cavity modes directly coupled to an ancilla for operations, and a storage unit with a larger number of modes for information storage. A coupler then enables SWAP-type operations between the two units, facilitating state transfer between storage and manipulator modes. In this way, the architecture potentially allows for hardware-efficient multimode control while minimizing crosstalk.
V.2 Generalization to other unselective cavity control protocols
Within the “all-to-one” multimode architecture, crosstalk error is ubiquitous across various cavity control protocols, though the degree of susceptibility varies. For protocols involving selective ancilla operations, such as the SNAP gate, crosstalk error is unavoidable due to the inherent selective nature of the protocol. This selectivity to photons in one cavity mode naturally makes it vulnerable to photons in other cavity modes. Conversely, there are protocols that rely on unselective ancilla operations, where ideally the ancilla operation does not depend on the photon number.
Generally, our technique to suppress crosstalk errors can be applied to multimode control protocols that involve unselective ancilla operations. In addition to the two-mode ECD protocol discussed in the main text, other protocols include, for example, the conditional not displacement gate proposed in Ref. 37. Furthermore, developing robust pulses for a coupled ancilla-cavity system could be extended to protocols with sideband drives [95, 86, 27]. By adapting our robust pulse development approach to these contexts, it is possible to mitigate crosstalk errors and enhance the fidelity of a broader range of quantum control protocols.
VI Conclusions
@@ Appendix A Multimode universal control with ECD protocol
Appendix B Detailed derivation of numerical simulation for a noisy driven coupled cavity-ancilla system
In this section, we provide a detailed step-by-step derivation of the procedure for numerically simulating a driven cavity-ancilla system in the presence of noise.
B.1 Dispersive Hamiltonian for the undriven system
The above approximations are valid in the regime where $|\Delta_{i}|\gg|K_{\text{q}}|$, consistent with the parameter regime considered throughout the paper. Specifically, we fix the parameter set: $K_{\text{q}}/2\pi=-200\,\text{MHz},\Delta_{1}/2\pi=\Delta_{2}/2\pi=2\,\text{GHz}$, and vary the coupling strengths $g_{i}$, taken to be identical for both cavity modes. Given that the dispersive shift $\chi_{i}$ is a directly measurable experimental quantity, we employ it to serve as the proxy for variations in the coupling strengths.
B.2 Rotating-frame Hamiltonian for the driven system
[…]
B.3 Displaced-frame Hamiltonian for the driven system
When dealing with weak cavity drives, the rotating-frame Hamiltonian can be used to conveniently simulate the system dynamics. However, for large drive amplitudes where the cavity is highly populated, this approach demands a large truncation level of the cavity Hilbert space to ensure numerical convergence, thereby increasing computational cost. A more computationally efficient strategy involves transforming the Hamiltonian to a drive-dependent displaced frame, in which the cavity displacement is eliminated (i.e., terms proportional to $\hat{a}_{i}$ or $\hat{a}_{i}^{\dagger}$ are removed, see details in Ref. 97). As a result, simulations in the displaced frame require a considerably smaller dimension for the truncated Hilbert space.
$\partial_{t}\alpha_{i}=i\chi_{i}\alpha_{i}/2-iK_{i}|\alpha_{i}|^{2}\alpha_{i}-iK_{12}\alpha_{i}|\alpha_{\bar{i}}|^{2}-i\Omega_{i}(t),$ (28)
$\hat{H}_{\text{diag}}(t)$ $=\sum_{i=1,2}\bigg{[}(2K_{i}|\alpha_{i}|^{2}+K_{12}|\alpha_{\bar{i}}|^{2})\hat{a}_{i}^{\dagger}\hat{a}_{i}$
[…]
B.4 Displaced-frame master equation for the driven open system
$\partial_{t}\eta_{i}=i\chi_{i}\eta_{i}/2-iK_{i}|\eta_{i}|^{2}\eta_{i}-iK_{12}\eta_{i}|\eta_{\bar{i}}|^{2}-i\Omega_{i}(t)-\kappa_{i}\eta_{i}/2,$
Appendix C Effects from sixth-order transmon nonlinearity
Here, we study how the sixth-order nonlinearity of the transmon affects the ECD protocol. Specifically, we consider the following transmon Hamiltonian:
$\hat{H}_{\text{trans}}=\omega_{\text{q}}\hat{q}^{\dagger}\hat{q}+\frac{K_{\text{q}}}{2}\hat{q}^{{\dagger}2}\hat{q}^{2}+\frac{G_{\text{q}}}{3}\hat{q}^{{\dagger}3}\hat{q}^{3},$ (33)
with the nonlinearities $K_{\text{q}}\approx-E_{\text{C}}$ and $G_{\text{q}}\approx\frac{E_{\text{C}}}{6}\sqrt{\frac{2E_{\text{C}}}{E_{\text{J}}}}$. In the typical transmon regime with $E_{\text{C}}\ll E_{\text{J}}$, we have $G_{\text{q}}\ll K_{\text{q}}$. When the transmon is dispersively coupled to a cavity, the sixth-order nonlinearity contributes in two ways: First, it adds minor corrections to the nonlinearities in Eq. (26). Second, it generates high-order cavity nonlinear terms such as $G_{i}\hat{a}_{i}^{{\dagger}3}\hat{a}^{3}_{i}/3$, with $G_{i}\approx(g_{i}/\Delta_{i})^{6}G_{\text{q}}$. When the cavity is strongly displaced in the ECD protocol, the above term might be non-negligible compared with the contribution from the cavity self-Kerr. This can be analyzed by generalizing Eq. (28), which determines the classical trajectory,
$\begin{split}\partial_{t}\alpha_{i}=&i\chi_{i}\alpha_{i}/2-iK_{i}|\alpha_{i}|^{2}\alpha_{i}-iG_{i}|\alpha_{i}|^{4}\alpha_{i}\\ &-iK_{12}\alpha_{i}|\alpha_{\bar{i}}|^{2}-i\Omega_{i}(t).\end{split}$ (34)
Therefore, the effect from the sixth-order nonlinearity is negligible when the following condition is satisfied:
$|\alpha|\ll\sqrt{\frac{K_{i}}{G_{i}}}\approx\sqrt{6}\left(\frac{g_{i}}{\Delta_{i}}\right)^{-1}\left(\frac{E_{\text{J}}}{2E_{\text{C}}}\right)^{1/4}.$ (35)
Within the transmon regime ($E_{\text{J}}\gg E_{\text{C}}$), the above bound is much larger than the one imposed by the critical photon number [36], with a ratio of $6(2E_{\text{J}}/E_{\text{C}})^{1/4}$. This suggests that the contribution from sixth-order transmon nonlinearity is negligible as long as the average photon in the cavity is below the critical photon number.
Appendix D Ancilla gate decomposition for efficient experimental calibration